OpenStudy (anonymous):

Which equation is correct? Select one: a. sec x° = opposite ÷ adjacent b. cosec x° = opposite ÷ adjacent c. cosec x° = hypotenuse ÷ opposite d. sec x° = adjacent ÷ hypotenuse

OpenStudy (anonymous):

@DK94

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

cosec x =1/(sin x), & since sin x is equal to opposite/hypotenuse the reciprocal of that would be c. hypotenuse/opposite

OpenStudy (anonymous):

Oh okay that makes sense, thanks love :3

OpenStudy (anonymous):

\[cosec(x)=\frac{1}{\sin(x)}\]

OpenStudy (anonymous):

\[\sin(x)=\frac{opposite}{hypotenuse}\] oh, haha. guess u got it. welcome miss @savvvvannahh